1. DeWitt, Corp just paid a dividend of $1 and the dividend will be growing at a constant rate of 18% for 2 years, and after that it will be growing at 3%. What is the intrinsic value of DeWitt's stock if investors require a rate of return of 7.0%? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

2.

A firm that has an ROE of 12% is considering increasing its dividend payout. The stockholders of the firm desire a dividend yield of 4% and a capital gain yield of 9%. Given this information, which of the following statements is (are) correct?

I. All else equal, the firm's growth rate will accelerate after the payout change.
II. All else equal, the firm's stock price will go up after the payout change.
III. All else equal, the firm's P/E ratio will increase after the payout change.

IV. None of the above.

Multiple Choice

  • I and II only

  • I only

  • II and III only

  • IV only
Answer #1

Hi,

As per policy we will solve only first question only.

1) Here Current dividend D0 = $1

Dividend growth rate for 2 years = 18%

Hence year 1 dividend D1 = 1*(1+18%) = $1.18

Present Value of year 1 dividend = 1.18/(1+7%) = $1.10

Year 2 dividend = D2 = 1.18*(1+18%) = $1.3924

Present Value of dividend = 1.3924/(1+7%)^2 = $1.22

Constant growth rate = 3%

So terminal value = D2*(1+g)/( k-g)

= 1.3924*(1+3%)/(7%-3%)

= $35.85

Present Value of terminal = 35.85/(1+7%)^2

= $31.32

Intrinsic value of stock will be sum of present value of future cash flow.

Intrinsic Value of Stock = 1.1 + 1.22 + 31.32

= $33.64

Thanks